OpenStudy (anonymous):

What is the slope of the tangent line to the graph of h(t) at t=2/3π if h(t)=cot(t)

OpenStudy (nenadmatematika):

first: do you mean h(t)=cos(t) and is t=2/(3pi) or t=(2/3)pi....?

OpenStudy (anonymous):

t=(2/3)pi

OpenStudy (accessdenied):

1) Find the derivative of h(t) 2) Plug in the value of t into derivative.

OpenStudy (nenadmatematika):

yes :D

OpenStudy (anonymous):